OpenStudy (anonymous):

what is the axis of symetry in of the parabola for y=2x^2+1

OpenStudy (anonymous):

y'=4x=0, x=0, which is the axis of symmetry.

OpenStudy (anonymous):

Also note we have no horizontal shifts, just a vertical one. Thus it defaults to x=0.

OpenStudy (anonymous):

dont u use -b/2a and get 1/4?

Directrix (directrix):

b = 0

OpenStudy (anonymous):

Ninja'd by Directrix. What he said.

Directrix (directrix):

y=2x^2+1 y = 2 x^2 + 0x + 1 a = 2, b = 0, and c = 1

OpenStudy (sasogeek):

i didn't think you could use differentiation to find the axis of symmetry... good to know :)

OpenStudy (anonymous):

That's how you derive -b/2a, by differentiation.

OpenStudy (sasogeek):

i see

OpenStudy (anonymous):

?whaaaa are u sure cause when i use that formula i get -1/4! btw im in 8th grade so please explain easily

Directrix (directrix):

y=2x^2+1 y=2x^2+1 y = 2 x^2 + 0x + 1 a = 2, b = 0, and c = 1 x-coordinate of vertex of parabola and also through which the axis of symmetry passes. x = -b/(2a) x = -(0)/ [(2)(2)] x = -0 / 4 x = 0 x=0 is the equation of the line of symmetry; it is also the y-axis. Take a look at the graph of the parabola and the axis of symmetry at the link: http://www.wolframalpha.com/input/?i=%28y%3D2x%5E2%2B1%29+and+%28x%3D0%29 A person is never too young to begin learning more mathematics nor too old.